Explain: The physical meaning of an uncertainty product. (You may use anyone discussed in the class.)

Explain:

The physical meaning of an uncertainty product. (You may use anyone discussed in the class.)
Answer #1

(\Delta x)(\Delta p_{x})\geq \frac{h}{4\pi }

\Delta x is uncertainty in position and \Delta p_{x} is uncertainty in momentum. Physically it describe that both position and momentum both can not be measure simultaneously. If position is measure accurately it will increase uncertainty in momentum. This is the basic rule of nature not an experimental error. Uncertainty principle is basics for quantum mechanics.

I will be glad to see your comment if you have any query and thumb up if you are satisfied. Thanks...